OpenStudy (anonymous):

Simplify the following expression: a + 2c - 5b - c + a - b 2a - 6b + 2c 2a - 6b + c a - 6b + c a - 5b + 2c

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

a + 2c - 5b - c + a - b = a + a - 5b - b + 2c - c = 2a - 6b + c

OpenStudy (anonymous):

do you understand?

OpenStudy (skullpatrol):

@ryanvarghese12790 could please provide more explanation?

OpenStudy (anonymous):

so answer is 2a-6b+c

OpenStudy (anonymous):

do you follow me?

OpenStudy (anonymous):

skullpatrol?

OpenStudy (skullpatrol):

How did you make subtraction commutative?

OpenStudy (anonymous):

It is easy to understand You put all the terms together and simply work out: (1)a+(1)a =2a then, -5b--(1)b = 6b then, 2c-(1)c = (1)c

OpenStudy (anonymous):

do you get it skullpatrol

OpenStudy (skullpatrol):

How did you make subtraction commutative?

OpenStudy (skullpatrol):

\[a -b \neq b -a\]

OpenStudy (anonymous):

the answer would be 2a+c−6b

OpenStudy (anonymous):

yes